# Tilting representations of finite groups of Lie type

### v1

@@
Abstract. Let $\mathbf{G}$ be a connected reductive group over a finite field $\mathbb{F}_{q}$ of characteristic $p>0$. In this paper, we study a category which we call Deligne–Lusztig category $\mathcal{O}$ and whose definition is similar to category $\mathcal{O}$. We use this to construct a collection of representations of $\mathbf{G}(\mathbb{F}_{q})$ which we call the tilting representations. They form a generating collection of integral projective representations of $\mathbf{G}(\mathbb{F}_{q})$. Finally we compute the character of these representations and relate their expression to previous calculations of Lusztig and we then use this to establish a conjecture of Dudas–Malle.
@@ 1. Introduction
Let $p>0$ be a prime and $q=p^{r}$ a power of $p$ and let $\mathbb{F}_{q}\subset\overline{\mathbb{F}}_{q}=k$ be the finite field with $q$ elements and its algebraic closure. Let $\mathbf{G}$ be a connected reductive group over $k$ equipped with a Frobenius endomorphism $\mathrm{F}:\mathbf{G}\to\mathbf{G}$, i.e. a purely inseparable isogeny such that a power of $\mathrm{F}$ is the Frobenius coming from an $\mathbb{F}_{q^{s}}$-structure. We fix $\mathbf{B}=\mathbf{T}\mathbf{U}$ a Borel pair that is stable under $\mathrm{F}$ and we denote by $\mathbf{W}$ the corresponding Weyl group. There are two goals to this paper
@@ 1.1.1. Category $\mathcal{O}$
as the category of $\ell$-adic sheaves, where $\ell\neq p$ is a prime number, on the flag variety of $\mathbf{G}$ that are locally constant along $\mathbf{U}$-orbits. Its perverse heart is an abelian category which has the structure of a highest weight category [BGS96] and therefore has very good homological properties. While it is known that this category provides a categorification of the Hecke algebra, it is not equipped with a monoidal structure. To witness the algebra structure of the Hecke algebra, one has to pass to Hecke categories. They come in two flavors.
@@
    the object $A$ is tilting if is has both a $\Delta$ and a $\nabla$-filtration.
@@
1.1.2. Deligne-Lusztig catgory $\mathcal{O}$
@@
which we call the horocycle stack. We had previously introduced this stack in [Ete24] [Ete24]. We define
@@
the category of $\Lambda$-sheaves on it where $\Lambda\in\{\overline{\mathbb{F}}_{\ell},\overline{\mathbb{Q}}_{\ell},\overline{\mathbb{Z}}_{\ell}\}$. We call this category ‘Deligne-Lusztig category $\mathcal{O}$’. We consider it as a variant of the classical category $\mathcal{O}$. To put some emphasis on the comparison, category $\mathcal{O}$ is classically realized as the category of sheaves on
@@
    The category $\mathcal{O}^{\mathrm{DL}}$ is compactly generated and the standard and costandard sheaves are compact.
@@
    For all pairs $(w,\chi)$, where $w\in\mathbf{W}$ and $\chi$ is a character of ${\mathbf{T}}^{w\mathrm{F}}$, there exists a unique (up to isomorphism) indecomposable tilting sheaf $T_{w,\chi}$ supported on the closure of $\mathbf{B}w\mathbf{B}$ containing $\Delta_{w,\chi}$ with multiplicity $1$ in a $\Delta$-filtration, moreover any indecompsable tilting sheaf is of this form.
@@
    We compare $\mathcal{O}^{\mathrm{DL}}$ with the Hecke categories $\mathcal{H}^{\mathrm{eq}}$ and $\mathcal{H}^{\mathrm{mon}}$ and their twisted variants of [LY20] and [Gou22] in section 2.2.
@@
It follows from this discussion that the structure of $\mathcal{O}^{\mathrm{DL}}$ is essentially controlled by Kazhdan-Lusztig type combinatorics.
@@ 1.2. Tilting representations
1.2.1. Deligne-Lusztig theory
Classically, the representation theory of ${\mathbf{G}}^{\mathrm{F}}$ is studied using Deligne-Lusztig theory [DL76]. One first defines two collections of varieties
@@ 1.2.1. Deligne-Lusztig theory
$[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]=\sum_{i}(-1)^{i}[\mathrm{IH}^{i}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$
of the intersection cohomology of the closure of the variety $X(w)$ in $\mathbf{G}/\mathbf{B}$ (which is usually a singular variety). There is an explicit relation between this character and the character $R_{w}^{1}$ expressed in terms of Kazhdan-Lusztig polynomials done in [Lus84], this relation is a reflection of some geometric relation between Deligne-Lusztig theory and the Hecke category.
@@ 1.2.2. Tilting representations
[…]
@@
    Tilting representations are perfect complexes.
  3. $(iii)$.
    If $q$ is large enough, then tilting representations are concentrated in degree 00 and are projective objects.
@@ 1.2.3. A conjecture of Dudas–Malle
More generally, Lusztig, in [Lus84], has given a formula for the Alvis-Curtis dual of $[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$. We reinterpret his calculation as follows
@@
The calculation of Lusztig and our theorem admit refinements which take into account the weights of the Frobenius endomorphism acting on the cohomology (and interesection homology) of the Deligne-Lusztig varieties. This calculation seems to indicate some deep interplay between Koszul duality for category $\mathcal{O}$ and Alvis-Curtis duality.
Finally, let $\delta>0$ be the smallest integer such that $(\mathbf{G},\mathrm{F}^{\delta})$ is split. Then $\mathrm{F}^{\delta}$ acts on the cohomology of $[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$. Let $\lambda\in\overline{\mathbb{Q}}_{\ell}$ and denote by $[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\lambda]]$ the character of the generalized eigenspace of $\mathrm{F}^{\delta}$ on $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})$ for the eigenvalue $\lambda$. Let $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$ and let $[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\overline{\lambda}]]$ be the character of
$[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\overline{\lambda}]]=\bigoplus_{\lambda}[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\lambda]]$
@@
Assume that $q$ is large enough (as in Theorem 1.2.4). There exists $\ell_{0}$ depending only on $\mathbf{G}$ (not on $\mathrm{F}$) such that for all $\ell\geq\ell_{0}$, $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$, $w\in\mathbf{W}$ the character of ${\mathbf{G}}^{\mathrm{F}}$
@@ 1.3. Notation and conventions
the ind-completion of the usual derived category of constructible sheaves on $X$. If the coefficient ring is clear, we will write $\mathrm{D}(X,\Lambda)=\mathrm{D}(X)$. Let $f:X\to Y$ be a morphism of stacks, whenever they are defined, we will denote by $f_{*},f_{!},f^{*},f^{!}$ the usual pushforward and pullbacks. We fix a square root of $p$ in $\overline{\mathbb{Z}}_{\ell}$. Whenever $X$ is a stack defined over $\mathbb{F}_{q}$, this defines a half Tate twist $(\frac{1}{2})$ on sheaves on $X_{\overline{\mathbb{F}}_{q}}$ with a Weil-structure.
We denote by $\mathrm{Hom}(-,-)$ the functor denoted by $R\mathrm{Hom}$ and we denote by $\mathrm{Hom}^{i}$ its $i$-th cohomology groups and by $\mathrm{End}(X)$ the functor $\mathrm{Hom}(X,X)$ (where everything is derived).
@@ 1.4. Acknowledgments
The author thanks Olivier Dudas for suggesting the problem and for comments on a draft of this paper. We thank Jean-François Dat and Olivier Dudas for their support during the preparation of this paper. We thank Gunter Malle for doing a careful reading of a draft of this paper. This paper was written while the author was a guest at the Max Planck Institute for mathematics in Bonn. The author thanks Cédric Bonnafé and Gerhard Hiss for helpful discussions.
2. Deligne-Lusztig category $\mathcal{O}$
@@
which we call the horocycle stack. We denote by $\mathcal{O}^{\mathrm{DL}}_{\Lambda}$ the category $\mathrm{D}(\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}},\Lambda)$. We call this category the Deligne-Lusztig category $\mathcal{O}$. The goal of this section is to discuss the homological structure of this category. When the coefficients are clear, we will drop the index $\Lambda$ from the notation. We denote by $\mathcal{O}_{\mathrm{c}}^{\mathrm{DL}}$ the full subcategory of constructible sheaves.
@@ 2.1. Stratification
The category $\mathcal{O}^{\mathrm{DL}}$ is compactly generated. The compact objects are exactly the sheaves that are constructible and compact when *-pullbacked (equivalently !!!-pullbacked) to all strata. In particular, the standard and costandard objects are compact.
@@
The category $\mathcal{O}^{\mathrm{DL}}$ is is generated by the objects $\Delta_{w,\chi}$. Since the immersions $j_{w}$ are of finite presentation both functors $j_{w}^{!}$ and $j_{w,*}$ are continuous hence their left adjoints preserve compact objects. If follows that the standard sheaves form a set of compact generators of the category.
@@ 2.3. Tilting objects
For the characterization of indecomposable tilting objects, we first note that the functor $\mathfrak{p}_{!}[\dim\mathbf{T}]$ sends tilting objects onto tilting objects by Lemma 2.2.1. For free monodromic Hecke categories, the corresponding statement is established in [BY13] in the unipotent case and [Gou22] in the non-unipotent case. This guaranties the existence of a tilting sheaf supported on the closure of $\mathbf{B}w\mathbf{B}$ and such that the multiplicity of $\Delta_{w,\chi}$ in one of its $\Delta$-flags is $1$. Since the category of tilting sheaves is Krull-Schmidt there exists a direct summand that is indecomposable and satisfies both of these properties. The unicity can be proved as in [Ric16, Theorem 7.14]. ∎
@@
    $j_{v}^{*}B=j_{v}^{*}A$ if $v>_{\mathrm{tot}}w$ and is therefore perverse and projective,
@@
Let $T,T^{\prime}$ be two tilting sheaves then $\mathrm{Hom}(T,T^{\prime})$ is concentrated in degree 00.
@@
By the five lemma, this reduces to the fact that $\mathrm{Hom}(\Delta,\nabla)$ is concentrated in degree 00. ∎
@@
Proposition 2.3.8.
There is a unique weight structure on $\mathcal{O}^{\mathrm{DL},\omega}$ whose heart is $\mathcal{O}^{\mathrm{DL}}_{\mathrm{tilt}}$. Furthermore, the weight complex functor
$\mathcal{O}^{\mathrm{DL},\omega}\to\mathrm{K}^{b}(\mathcal{O}_{\mathrm{tilt}}^{\mathrm{DL}})$
is an equivalence.
Proof.
Since tilting sheaves generate the category by Proposition 2.3.7 and the mapping spaces between them are concentrated in degree 00 by Lemma 2.3.6, the same proof as in [EE24, Corollary 6.22] applies. ∎

### v2

@@
Abstract. Let $\mathbf{G}$ be a connected reductive group over a finite field $\mathbb{F}_{q}$ of characteristic $p>0$. In this paper, we study a category which we call Deligne–Lusztig category $\mathcal{O}$ and whose definition is similar to category $\mathcal{O}$. We use this to construct a collection of representations of $\mathbf{G}(\mathbb{F}_{q})$ which we call the tilting representations. They form a generating collection of integral projective representations of $\mathbf{G}(\mathbb{F}_{q})$. Finally we compute the character of these representations and relate their expression to previous calculations of Lusztig and we then use this to establish a conjecture of Dudas–Malle of 2014.
@@ 1. Introduction
Let $p>0$ be a prime and $q=p^{r}$ a power of $p$ and let $\mathbb{F}_{q}\subset\overline{\mathbb{F}}_{q}=k$ be the finite field with $q$ elements and its algebraic closure. Let $\mathbf{G}$ be a connected reductive group over $k$ equipped with a Frobenius endomorphism $\mathrm{F}:\mathbf{G}\to\mathbf{G}$, i.e. a purely inseparable isogeny such that a power of $\mathrm{F}$ is the Frobenius coming from an $\mathbb{F}_{q^{s}}$-structure. We fix $\mathbf{B}=\mathbf{T}\mathbf{U}$ a Borel pair that is stable under $\mathrm{F}$ and we denote by $\mathbf{W}$ the corresponding Weyl group. This paper has two main goals :
@@ 1.1.1. Category $\mathcal{O}$
as the category of $\ell$-adic sheaves, where $\ell\neq p$ is a prime number, on the flag variety of $\mathbf{G}$ that are locally constant along $\mathbf{U}$-orbits. Its perverse heart is an abelian category which has the structure of a highest weight category [BGS96] and therefore has very good homological properties. While it is known that this category provides a categorification of the Hecke algebra, it is not equipped with a monoidal structure. to reveal the algebra structure of the Hecke algebra, one must pass to Hecke categories. They come in two flavors.
@@
    the object $A$ is tilting if it has both a $\Delta$ and a $\nabla$-filtration.
@@
1.1.2. Deligne–Lusztig catgory $\mathcal{O}$
@@
which we call the horocycle stack. We had previously introduced this stack in [Ete24]. We define
@@
the category of $\Lambda$-sheaves on it where $\Lambda\in\{\overline{\mathbb{F}}_{\ell},\overline{\mathbb{Q}}_{\ell},\overline{\mathbb{Z}}_{\ell}\}$. We call this category ‘Deligne–Lusztig category $\mathcal{O}$’. We consider it as a variant of the classical category $\mathcal{O}$. To put some emphasis on the comparison, category $\mathcal{O}$ is classically realized as the category of sheaves on
@@
    The category $\mathcal{O}^{\mathrm{DL}}$ is compactly generated, and the standard and costandard sheaves are compact.
@@
    For all pairs $(w,\chi)$, where $w\in\mathbf{W}$ and $\chi$ is a character of ${\mathbf{T}}^{w\mathrm{F}}$, there exists a unique (up to isomorphism) indecomposable tilting sheaf $T_{w,\chi}$ supported on the closure of $\mathbf{B}w\mathbf{B}$ containing $\Delta_{w,\chi}$ with multiplicity $1$ in a $\Delta$-filtration, moreover any indecomposable tilting sheaf is of this form.
@@
    We compare $\mathcal{O}^{\mathrm{DL}}$ with the Hecke categories $\mathcal{H}^{\mathrm{eq}}$ and $\mathcal{H}^{\mathrm{mon}}$ and their twisted variants of [LY20] and [Gou22] in section $2.$
@@
It follows from this discussion that the structure of $\mathcal{O}^{\mathrm{DL}}$ is essentially controlled by Kazhdan–Lusztig type combinatorics.
@@ 1.2. Tilting representations
1.2.1. Deligne–Lusztig theory
Classically, the representation theory of ${\mathbf{G}}^{\mathrm{F}}$ is studied using Deligne–Lusztig theory [DL76]. One first defines two collections of varieties
@@ 1.2.1. Deligne–Lusztig theory
$\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]=\sum_{i}(-1)^{i}[\mathrm{IH}^{i}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$
of the intersection cohomology of the closure of the variety $X(w)$ in $\mathbf{G}/\mathbf{B}$ (which is usually a singular variety). There is an explicit relation between this character and the character $R_{w}^{1}$ expressed in terms of Kazhdan–Lusztig polynomials done in [Lus84], this relation is a reflection of some geometric relation between Deligne–Lusztig theory and the Hecke category.
@@ 1.2.2. Tilting representations
$\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{G}}\xleftarrow{\alpha}\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{B}}\xrightarrow{\beta}\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}.$
@@
The next theorem has also been discovered independently by Zhu [Zhu25].
@@
    Tilting representations are concentrated in degree 0 and are projective representations.
@@ 1.2.3. A conjecture of Dudas–Malle
More generally, Lusztig, in [Lus84], has given a formula for the Alvis–Curtis dual of $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$. We reinterpret his calculation as follows
@@
The calculations of Lusztig and our theorem admit refinements which take into account the weights of the Frobenius endomorphism acting on the cohomology (and intersection homology) of the Deligne–Lusztig varieties. This calculation suggests some deep interplay between Koszul duality for category $\mathcal{O}$ and Alvis–Curtis duality.
Finally, let $\delta>0$ be the smallest integer such that $(\mathbf{G},\mathrm{F}^{\delta})$ is split. Then $\mathrm{F}^{\delta}$ acts on the cohomology of $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$. Let $\lambda\in\overline{\mathbb{Q}}_{\ell}$ and denote by $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\lambda]]$ the character of the generalized eigenspace of $\mathrm{F}^{\delta}$ on $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})$ for the eigenvalue $\lambda$. Let $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$ and let $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\overline{\lambda}]]$ be the character of
$\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\overline{\lambda}]]=\bigoplus_{\lambda}[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\lambda]]$
@@
There exists $\ell_{0}$ depending only on $\mathbf{G}$ (not on $\mathrm{F}$) such that for all $\ell\geq\ell_{0}$, $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$, $w\in\mathbf{W}$ the character of ${\mathbf{G}}^{\mathrm{F}}$
@@
Remark 1.2.7.
The hypothesis on $\ell$ in the theorem is to ensure that Kazhdan–Lusztig polynomials and $\ell$-Kazhdan–Lusztig polynomials agree. This ensure that if $T^{\overline{\mathbb{Z}}_{\ell}}_{w}$ is a $\overline{\mathbb{Z}}_{\ell}$ tilting representation, then the unipotent part of $T^{\overline{\mathbb{Z}}_{\ell}}_{w}[\frac{1}{\ell}]$ is $T_{w}^{\overline{\mathbb{Q}}_{\ell}}$ whose character is known. In general, the unipotent part of $T^{\overline{\mathbb{Z}}_{\ell}}_{w}[\frac{1}{\ell}]$ decomposes as
$T^{\overline{\mathbb{Q}}_{\ell}}_{w}\oplus_{v<w}T_{v}^{\overline{\mathbb{Q}}_{\ell},\oplus n_{w,v}}$
where $n_{w,v}\in\mathbb{N}$ are certain multiplicities.
@@ 1.3. Notation and conventions
the ind-completion of the usual derived category of constructible sheaves on $X$. If the coefficient ring is clear, we will write $\mathrm{D}(X,\Lambda)=\mathrm{D}(X)$. Let $f:X\to Y$ be a morphism of stacks, whenever they are defined, we will denote by $f_{*},f_{!},f^{*},f^{!}$ the usual pushforward and pullbacks. We fix a square root of $p$ in $\overline{\mathbb{Z}}_{\ell}$. Whenever $X$ is a stack defined over $\mathbb{F}_{q}$, this defines a half Tate twist $(\frac{1}{2})$ on sheaves on $X_{\overline{\mathbb{F}}_{q}}$ with a Weil-structure. Given a stack $X$ with a Frobenius $\mathrm{F}:X\to X$, we denote by $\mathrm{D}^{b}_{c}(X)^{\mathrm{Weil}}$ the category of Weil sheaves on $X$, i.e. the category of pairs $(A,\phi)$ where $A\in\mathrm{D}^{b}_{c}(X)$ and $\phi:\mathrm{F}^{*}A\xrightarrow{\sim}A$.
We denote by $\mathrm{Hom}(-,-)$ the derived functor of endomorphism and we denote by $\mathrm{Hom}^{i}$ its $i$-th cohomology groups and by $\mathrm{End}(X)$ the functor $\mathrm{Hom}(X,X)$ (where everything is derived).
@@ 1.4. Acknowledgments
The author thanks Olivier Dudas for suggesting the problem and for comments on a draft of this paper. We thank Jean-François Dat and Olivier Dudas for their support during the preparation of this paper. We thank Gunter Malle for doing a careful reading of a draft of this paper. This paper was written while the author was a guest at the Max Planck Institute for mathematics in Bonn. The author thanks Cédric Bonnafé and Gerhard Hiss for helpful discussions. The author thanks Xinwen Zhu for comments on a previous version of this paper.
2. Deligne–Lusztig category $\mathcal{O}$
@@
which we call the horocycle stack. We denote by $\mathcal{O}^{\mathrm{DL}}_{\Lambda}$ the category $\mathrm{D}(\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}},\Lambda)$. We call this category the Deligne–Lusztig category $\mathcal{O}$. The goal of this section is to discuss the homological structure of this category. When the coefficients are clear, we will drop the index $\Lambda$ from the notation. We denote by $\mathcal{O}_{\mathrm{c}}^{\mathrm{DL}}$ the full subcategory of constructible sheaves.
@@ 2.1. Stratification
The category $\mathcal{O}^{\mathrm{DL}}$ is compactly generated. The compact objects are exactly the sheaves that are constructible and compact when *-pullbacked (equivalently !!-pullbacked) to all strata. In particular, the standard and costandard objects are compact.
@@
The category $\mathcal{O}^{\mathrm{DL}}$ is generated by the objects $\Delta_{w,\chi}$. Since the immersions $j_{w}$ are of finite presentation both functors $j_{w}^{!}$ and $j_{w,*}$ are continuous hence their left adjoints preserve compact objects. If follows that the standard sheaves form a set of compact generators of the category.
@@ 2.3. Tilting objects
For the characterization of indecomposable tilting objects, we first note that the functor $\mathfrak{p}_{!}[\dim\mathbf{T}]$ sends tilting objects to tilting objects by Lemma 2.2.1. For free monodromic Hecke categories, the corresponding statement is established in [BY13] in the unipotent case and [Gou22] in the non-unipotent case. This guarantees the existence of a tilting sheaf supported on the closure of $\mathbf{B}w\mathbf{B}$ and such that the multiplicity of $\Delta_{w,\chi}$ in one of its $\Delta$-flags is $1$. Since the category of tilting sheaves is Krull-Schmidt there exists a direct summand that is indecomposable and satisfies both of these properties. The uniqueness can be proved as in [Ric16, Theorem 7.14]. ∎
@@
    $j_{v}^{*}B=j_{v}^{*}A$ if $v>_{\mathrm{tot}}w$ and is perverse and projective by induction,
@@
Let $T,T^{\prime}$ be two tilting sheaves then $\mathrm{Hom}(T,T^{\prime})$ is concentrated in degree 0.
@@
By the five lemma, this reduces to the fact that $\mathrm{Hom}(\Delta,\nabla)$ is concentrated in degree 0. ∎